OpenStudy (anonymous):

A sector of a circle with a 120 degree arc has an area of 75(pi)in^2. Find the radius.

OpenStudy (anonymous):

Basically, you must first find out the formula for the sector of a circle. From there on, you can find the radius. The formula of a sector is: \[\theta*r ^{2}\] The reason as to why this is the formula is simple: because a sector is actually a fraction of the full circle. Now you can easily find the radius of that circle, if you set up an equation: \[\theta r^{2} = area\] and solve for r

OpenStudy (anonymous):

Maybe I didn't explain the reason as to why that is the formula properly. As I was saying, a sector is just a part of the circle. What part? well, it's \[\frac{\theta}{2 \pi}\] parts of the circle. So you do \[\frac{\theta}{2 \pi} * \pi r ^{2} = \frac{\theta r ^{2}}{2}\] Thus, it looks like formula I gave you in the first post was wrong.

OpenStudy (anonymous):

you can see that 120/360 is a third.....

OpenStudy (anonymous):

my intention was to give a general approach for solving similar problems